Paul is 30 years older than his son Ron. Five years ago he was four times as old as his son was. How many years old is Ron now?
Answer by lwsshak3(11628)

Paul is 30 years older than his son Ron. Five years ago he was four times as old as his son was. How many years old is Ron now?
***
let x=Ron's age now
x+30=Paul's age now
..
5 yrs ago:
x-5=Ron's age
x+30-5=x+25
..
x+25=4(x-5)
x+25=4x-20
3x=45
x=15
Ron's age now=15